OpenStudy (anonymous):

Medal for the best answer!!!!!!!! Write an equation of the line, in slope-intercept form, that passes through the point (0, 4) and has the slope, m = 2.

OpenStudy (anonymous):

\(y=mx+b\) has slope \(m\) and \(y\) intercept \(b\) you are given both of those numbers

OpenStudy (anonymous):

so i already have the answer

OpenStudy (anonymous):

yes, pretty much you have nothing to find, plug the numbers in

OpenStudy (anonymous):

so would it be 0=2x+4

OpenStudy (anonymous):

no replace the \(0\) by \(y\) it is \(y=mx+b\) not \(0=mx+b\)

OpenStudy (anonymous):

so then whats the answer?